OpenStudy (anonymous):

What is 840,000 in scientific notation? a. 840 x 10^5 b. 84 x 10^-5 c. 8.4 x 10^5 d. 8.4 x 10^-5

OpenStudy (anonymous):

C is the correct notation

OpenStudy (anonymous):

For each 0 Forward sue from a power for 10 . Look at this example : 10=10^1____________ number of 0 = 1 100=10^2 1000=10^3 and ... So : 600000000=6 * 10^8 Got it ?;)

OpenStudy (anonymous):

@cj392691

OpenStudy (unklerhaukus):

scientific notation has only one digit before the decimal point,